OpenStudy (cccmpt2):

Find the y intercept: 2x + y = 7 The answer is y = 7/2, but I don't know how to get it.

OpenStudy (cccmpt2):

that is what my study guide says...?

OpenStudy (cccmpt2):

i was able to calculate 2x+3y=10 and got y=10/3 and the answers matched the guide...

Directrix (directrix):

Did you write this question correctly: --> Find the y intercept:

OpenStudy (cccmpt2):

i don't know if that helps any???

OpenStudy (cccmpt2):

aha, thank you Directrix! it is late and i am not reading well...solve for X!!!!! i think i should call it a night here soon!

Directrix (directrix):

I don't know what this means: "calculate 2x+3y=10" ----------------- Please post your work. Something is not right here. It could be the question, your answer, a bad answer in the study guide, or something else.

OpenStudy (cccmpt2):

assuming x is 0... 2(0)+3y=10 3y=10 divide both sides by 3 y=10/3

Directrix (directrix):

The x-intercept lies on the x-axis. All points on the x-axis have a y-value of 0. So substitute 0 for y and solve for x.

OpenStudy (cccmpt2):

x=5

Directrix (directrix):

2x + y = 7 Find the x-intercept. Let y = 0 2x + 0 = 7 --> Solve this for x -->@cccmpt2

OpenStudy (cccmpt2):

sorry, i though you were talking about the other prob - 2x+3y=10 ;)

OpenStudy (cccmpt2):

so x = 7/2, just like the study guide answer showed. thank you for helping me! it was really simple, once i read the problem right LOL!

Directrix (directrix):

I only saw one problem in this thread. Are two problems here? If so, it is best to put one problem per thread because if the thread runs long, it is easy to get mixed up as to which problem each other is talking about.

Directrix (directrix):

To what problem is this an answer--> x=5

Directrix (directrix):

The x-intercept of - 2x+3y=10 is not 5. I think you have a sign error.